$SPAD/src/input images8.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- Color gallery page 8
-- Sierpinsky's Tetrahedron
)read tetra
drawPyramid 4
-- Antoine's Rings
)read antoine
drawRings 2
-- Scherk's Minimal surface
)read scherk
drawScherk(3,3)
-- Ribbon Plot of [x**i for i in 1..5]
)r ribbons
)set message test off
drawRibbons([x**i for i in 1..5], x=-1..1, y=0..2)
)set message test on

———-

References^{[1] nothing}

4